OpenStudy (anonymous):

Which word phrase represents the variable expression? 5n + 8 A. five more than eight times a number B. eight more than five times a number C. five times the sum of a number and eight D. eight more than the sum of a number and five

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

I THINK IT IS C

OpenStudy (anonymous):

Right, let's start removing the one we think don't belong

OpenStudy (anonymous):

D & B

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

BECAUSE THEIR IS NO 8

OpenStudy (anonymous):

FROM THE BEGINNING

OpenStudy (anonymous):

Read "more than" as "+", and "times" as "*"

OpenStudy (anonymous):

hm, let's start making the sentences in to equations: A. five more than eight times a number 5+8n

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Ok, C. five times the sum of a number and eight 5(n+8)

OpenStudy (anonymous):

YES I THINK IT IS C

OpenStudy (anonymous):

ok: D. eight more than the sum of a number and five 8(n+5)

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

YOU THERE

OpenStudy (anonymous):

sorry, meant 8+n+5. It's says 8 more than the sum of a number which means 8+n right. Then you have "and five" which is +5

OpenStudy (anonymous):

anyway, B. eight more than five times a number 8+5n

OpenStudy (anonymous):

IS THAT THE ANSWER OR NO

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

yes it is. it's B

OpenStudy (anonymous):

TWO MORE

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which is an equation for the given word problem? After Nancy deposits her birthday gift of $20.00 into her savings account, her balance is $225.50. Which equation can be used to find the amount, n, in her account before she made the deposit? A. 20.00 – n = 225.50 B. n + 225.50 = 20.00 C. n – 225.50 = 20.00 D. 20.00 + n = 225.50

OpenStudy (anonymous):

What do ya think

OpenStudy (anonymous):

IM GOING WITH D. BECAUSE SHE HAD $225.50 SHE DEPOSIT $20.00 SO IT WOULD BE 20.00 + n = 225.50

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

One More Question

OpenStudy (anonymous):

K

OpenStudy (anonymous):

One More Question

OpenStudy (anonymous):

One More Question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@ganeshie8 can you close the question?

OpenStudy (anonymous):

When it says 5n+8 the + signs is read as "more than" so if I know correctly, you have to say the equation backwards, if it says 8 more than 5n, it is 5n+8 but if it is 5n less than 8 then it is 8-5n, get it? its not a joke.

OpenStudy (anonymous):

@Lovelarap you should be a little nicer.

OpenStudy (anonymous):

sorry....

OpenStudy (anonymous):

That's okay.